WAC Notes December 15 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes CREX << Instructions

WAC Notes December 14 2019 << >> WAC Notes January 8 2020

December 15 2019

Daily Meeting BlueJeans

Production Run List - Slug 102 - IHWP OUT - Flip RIGHT - Production

  • Good - Run 5454 - Ca48 150uA - good elastic peak on maindet alignment now, 776 ppm combined asym width
  • Good - Run 5455
  • Good - Run 5456 - Definitely has a 0.3 mm BPM 12X shift that could be a problem for elastic peak separation - 400k onward cut due to excursions
  • Good - Run 5457
  • Good - Run 5458 - Definitely has a 1.0 mm BPM 12X drift that could be a problem for elastic peak separation and slope finding
  • Good - Run 5459 - Definitely has a 1.0 mm BPM 12X drift that could be a problem for elastic peak separation and slope finding
  • Good - Run 5460 - Added a BPM 4eX > 0.3mm cut to remove a small excursion
  • Good - Run 5461

Production Run List - Slug 103 - IHWP IN - Flip RIGHT - Production

  • Good - Run 5462
  • Good - Run 5463 - still has large energy drifts
  • Good - Run 5464 - Added a BPM 4eX > 0.3mm cut to remove a small excursion
  • Good - Run 5465 - still has large energy drifts
  • Good - Run 5466
  • Good - Run 5467 - Large late BPM 12 Y shift, but not fatal on 4eY
  • Good - Run 5468 - Cutting out 20k-120k due to unstable and trippy beam. Still has large energy drifts
  • Good - Run 5469

Production Run List - Slug 104 - IHWP OUT - Flip RIGHT - Production

  • Good - Run 5471 - still has large energy drifts
  • Good - Run 5472 - Added a BPM 4eX > 0.3mm cut to remove a small excursion. cutting events 417000-421000, otherwise good
  • Good - Run 5473 - still has large energy drifts
  • Good - Run 5474 - Pedestal run, has a minirun of good data in it
  • Good - Run 5475 - FFB finally turned on
  • Good - Run 5476
  • Good - Run 5477 - Ca48 150uA - Cut out the first 45k events due to strong beam instability
  • Good - Run 5478
  • Good - Run 5486 - Really large motion in BPMs for a restoration run... maybe needs a cut later
  • Good - Run 5488
  • Good - Run 5489 - super trippy

Non-Production Runs and Bad Runs

  • Short junk - Run 5470 - Short, bad Aq, only BMOD data

Pedestal/Calibrations Updates

Issues

  • Slug 102-104 have FFB issues that prevent BPM 12X from staying still - many energy drifts over this slug which will be hard to fix
  • Slug 104 has FFB turn on mid way through the slug - this makes the optics/energy in the HRS more comfortable and stable
  • SAM 7 is updated and ATs are moved at the end of slug 103 and ATs are moved at the end of slug 104, so these are bifurcations
  • Fixed the prompt summary to report correct number of patterns: https://logbooks.jlab.org/entry/3753878
  • 12X energy shifting
    • Dispersion in the arc is measured as 0.4 MeV/mm
    • Dispersion in the HRS is 6.2 mm/MeV
    • Dithering does a 0.5 mm p2p fluctuation in 12 X -> 1.25 mm shift in the HRS Quartz
    • Normal running without FFB does 1 mm p2p fluctuations (over hour time scale) -> 2.5 mm shift in the HRS quartz
  • 5477.0 has a significant problem in JAPAN
  • Long vs. short bursts will require JAPAN rNd from Paul, etc.
  • The AT combos of interest are Ldd avg Rdd (for vertical transverse), Indd avg Outdd (most important for horizontal transverse), and dd's therein

BCMs

  • Run 5474 is a current ramp for calibrating BCM Cav 4CQ, AT Rs, and SAM 7

Pedestals and Calibrations (respin first 4 slugs)

History of calibrations spreadsheet

Committed: https://github.com/JeffersonLab/japan/commit/ddb956334b38acb124f67b1e841b5a02c761cc97

Other Status Updates